OpenStudy (anonymous):

If the coefficient is negative, you can factor out -1 (negative one) first. Then you can factor the rest of the expression more easily. After factoring the polynomial, how do you determine which values of the variable make the original expression undefined? Give an example of a polynomial expression, that has at least one undefined value and show it in factored form.

OpenStudy (ganpat):

yes, u can do that.. and do not forget to multiply by - sign at end..

OpenStudy (anonymous):

Do you have an example of a polynomial expression?

OpenStudy (ganpat):

well, i dont have any text book with me.. so let me think.. In between, lets c, if some one else comes up with answer.. @waterineyes , @ParthKohli

OpenStudy (anonymous):

I am not getting the question @Ganpat ??

OpenStudy (anonymous):

All I need is a polynomial expression that is in factored form. And that has an undefined value....

OpenStudy (ganpat):

say for example.. -2x2 - 2x = 0 so solving it.. - (2x2 + 2) = 0 2x2 + 2 = 0 2x (x +1) = 0 2x=0 or x+ 1 = 0 x = 0 or x = -1 is this what u r looking for ??????